OpenStudy (anonymous):

Find the decomposition of a(t) into tangenital and normal components at the point indicated

OpenStudy (anonymous):

\[r(t)=<e^t,1-t>\]\[t=0\]

OpenStudy (anonymous):

Okay, let's compute our derivatives:$$\mathbf{r}(t)=(e^t,1-t)\\\mathbf{r}'(t)=(e^t,-1)\\\mathbf{r}''(t)=(e^t,0)$$Can you determine our unit tangent vector from this?

OpenStudy (anonymous):

The second one!

OpenStudy (anonymous):

\[a_T = \frac{ r'(t)*r''(t) }{ ||r'(t)|| }\] \[a_N = \frac{ ||r'(t) \times r''(t)|| }{ ||r'(t)|| }\] * (dot product)

OpenStudy (anonymous):

@FutureMathProfessor that's not our unit tangent vector, that's just *a* tangent vector. To make it unit length, find its magnitude:$$\|\mathbf{r}'(t)\|=\sqrt{e^{2t}+1}$$and now we scale \(\mathbf{r}'\) by its reciprocal:$$\mathbf{T}(t)=\frac1{\sqrt{e^{2t}+1}}\mathbf{r}'(t)=\left(\frac{e^t}{\sqrt{e^{2t}+1}},-\frac1{\sqrt{e^{2t}+1}}\right)$$

OpenStudy (anonymous):

@Euler271 that is incorrect by the way :-)

OpenStudy (anonymous):

So then if I plugged in t=0, the vector would be (1/sqrt(2),-1/sqrt(2))

OpenStudy (anonymous):

You're finding the unit vector. I'm finding the actual magnitude (times the unit vector) Not sure what the question is asking

OpenStudy (anonymous):

@Euler271 that formula is incorrect for \(a_N\)... you need \(\|r''(t)\|\) in that denominator hehe.

OpenStudy (anonymous):

@FutureMathProfessor correct! Now, compute our acceleration vector at \(t=0\):$$\mathbf{r}''(0)=(1,0)$$Okay, now recall that for a twice-differentiable vector valued function \(\mathbf{r}'\) we know that our acceleration vector lies in the plane spanned by the unit tangent and unit normal vectors:$$$$